OpenStudy (bananas1234):

Which fraction shows 2.8 as the quotient of two integers?

zepdrix (zepdrix):

Hey Banana, do we have some options to choose from? :)

OpenStudy (bananas1234):

yea i can get them :)

OpenStudy (bananas1234):

14/5 12/5 9/4 9/2

zepdrix (zepdrix):

Remember that multiplying by 10 moves the decimal 1 place, so that will at least get us started with whole numbers.\[\Large\rm 2.8\left(\frac{10}{10}\right)=\frac{2.8\times10}{10}=\frac{28}{10}\]From there all we need to do is reduce this fraction to lower terms.

jimthompson5910 (jim_thompson5910):

hint: using a calculator or long division, you will find that 9/2 = 4.5 which is NOT the same as 2.8 allowing you eliminate choice D

zepdrix (zepdrix):

For 28 and 10, do they have FACTOR in common? Can we take something out of both of them to reduce the fraction?

OpenStudy (bananas1234):

Sorry i stepped away. I am back now. @zepdrix

OpenStudy (bananas1234):

would it be A? @zepdrix

OpenStudy (bananas1234):

Is A correct @jim_thompson5910

zepdrix (zepdrix):

A? Yayyy good job \c:/